### v1

@@ 3.2.2 A Dataset of Disaster App Reviews
Collecting Reviews: To answer RQ2 and RQ3, we needed user reviews of the 45 disaster apps. To gather user reviews of 45 disaster apps from the Google Play Store and App Store, we utilized an app-store-scraper  [^6] https://github.com/cowboy-bebug/app-store-scraper and google-play-scraper  [^7] https://github.com/JoMingyu/google-play-scraper. These tools require the app’s ID, name, and the specific country of the store. The app’s name is visible upon selection, whereas its ID and the relevant country code are identifiable through its URL. We collected 37,272 user reviews from 45 disaster apps.
Translation: The collected reviews encompassed a diverse linguistic range, particularly for apps from non-English speaking countries. This included apps such as ‘DEC112 2.0’, which has reviews in German, ‘Zachranka’ in Czech, and ‘Yurekuru Call’ in Japanese. Of 45 apps, 9 had reviews in languages other than English, totaling 2,793 reviews requiring translation. To address this, we used a Zero-Shot Translation model  [^8] https://huggingface.co/Helsinki-NLP to translate non-English reviews into English reviews.
@@
Exclusions: The next step was to exclude apps without user reviews and those with brief reviews comprising fewer than four words. The reason behind this exclusion is twofold: firstly, the absence of user reviews renders it impossible to address RQ2 and RQ3, which rely on user reviews. Secondly, short reviews, offering minimal commentary such as ‘good’, ‘not accurate’, and ‘works well’, provide insufficient information for meaningful analysis. Additionally, reviews with no rating were eliminated from the dataset, as they could not be used in the sentiment analysis stage. This filtering process reduced our app review dataset to comprising 28,161 reviews, laying a foundation for data analysis for RQ2 and RQ3.
@@ 4.1 RQ1: Key features of disaster apps
4.1.1 Approach
Analysis 1- App Features: To identify the available app features, we employed a three fold approach: reading the app descriptions, examining app image previews, and downloading the app for direct inspection. Some apps offered complete feature descriptions. Hence, we read such descriptions to identify and extract the app features. However, if an app lacked a clear description, we referred to the app image previews, typically a set of screenshots of the app contents in use displayed by the developer on the app’s page in the Play Store or App Store. If the feature images were incomplete, we downloaded the app and examined its features directly. After collecting data on each app’s features, we grouped them into three types of features based on four stages of the disaster life cycle: preparation, response, recovery and mitigation [7].
Analysis 2- Disaster Type Coverage: Using the same three fold approach described in the app feature analysis stage above, we also determined the types of disasters the app covers. This identified two types of apps: multi-hazard apps and single hazard apps. The single hazard app included four types of disasters: earthquake, wind, flood and fire.
Analysis 3- Geographic Coverage: To determine the area coverage of the app, we used the same three fold approach as before. Then, based on the area coverage, we grouped apps into two main categories: international and local. The local included three sub-categories: region-specific, national and territorial.
Analysis 4- Popularity and Quality Analysis: The popularity of an app can be gauged by its number of downloads, while user ratings offer insights into its perceived quality. These data can be obtained from the homepage of each app on the Play Store and App Store. For example, Figure 2 shows that the ‘Hurricane Tracker Pro’ app has been downloaded 5K+ and has a rate of 4.7.
Figure 2: App pricing scheme, number of downloads, and rating
Analysis 5- Monetization Strategies: Not all disaster support apps are made freely available by government agencies. To gather data about how app developers may be supported we examined the homepage of each app. The home page indicated whether the app contained advertising or offers in-app purchases (payments within the app to access certain features or remove advertisements). Additionally, on the install button, there was a price if the app is paid, and is labeled as ‘free’ if it is available without charge. Thus, each app can be completely free, or there can be a combination of profit-making methods. For example, as shown in Figure 2, users need to pay \$2.99 to download the ‘Hurricane Tracker Pro’ app. This app also offers a subscription feature. Another example is the ‘Wind Map Hurricane Tracker 3D’ app in Figure 2, which is free to download but contains advertising and offers subscriptions within the app.
4.1.2 Results
Analysis 1 – App Features: Each disaster support app provides one or more features related to natural disasters. The following is an explanation of the key disaster app features we categorized according to the stages of a disaster event: pre-disaster, disaster, and post-disaster. To view the number of apps with each feature type, refer to Table 1.
Table 1: App feature categorisation and number of apps
@@ 4.1.2 Results
Mitigation Help/Shelter Info 7
Recommendation Actions 2
@@
      Early Warning Alerts: These enable users to immediately locate a safe area and take action to protect themselves before a disaster occurs. Apps offering this feature are typically from Japan, especially those related to earthquakes. A good example is from ‘Yukureku Call’ app (see Figure LABEL:fig_ew).
@@
      Disaster Forecasting: This feature supports aspects of disaster and weather forecasting, which lasts longer than an early warning alert. While early warning alerts can be sent moments before a disaster, forecasting can be accessed anytime, even if a disaster does not occur. This feature is widely available in apps that monitor wind movements. For example, as seen in the ‘Seastorm Hurricane Tracker’ app (see the Figure LABEL:fig_forecast).
@@
      Real Time Alerts: This alert appears as some form of pop-up, informing the user of a disaster happening in the area according to their radius settings or specified area of interest. This feature differs from an early warning alert because notifications can arrive during a disaster, not beforehand. For instance, if a user sets a threshold for earthquakes measuring 3 on the Richter scale and above, they will receive a notification when such an event occurs.Such an example is provided in ‘QuakeFeed Earthquake Tracker’ app (see Figure LABEL:fig_at_disaster_1).
@@
    Figure 6: Features of Mitigation Stage
  • Features of Mitigation Stage: This stage is about taking steps to prevent the impact of future disasters.
    1. 1.
      Help/Shelter Info: In preparation for the next disaster, users can search for the nearest shelter using some apps. This will help them prepare evacuation routes and evacuation modes in preparation for a future disaster. Such an example if found in ‘Emergency: Severe Weather App’ (see Figure LABEL:fig_mit_disaster_1).
    2. 2.
      Recommendation Actions: This feature provides actions to take based on the post-disaster situation. Such recommendations for such actions can be found in ’NINA - Die Warn-App des BBK’ (see Figure LABEL:fig_mit_disaster_2). The following is a recommendation made by this app for the user, which is translated from German to English.
      Disaster: Due to a fire in a warehouse, there is heavy smoke development
      Recommended action: Please immediately close windows and doors. Turn off ventilation and air conditioning systems.”
Analysis 2 – Disaster Type Coverage: Based on the types of disasters covered, the apps can be divided into two: multi-hazard apps and single hazard apps. The single hazard app category (24 apps) includes four types of disasters: Earthquake, Wind-Related Disaster, Flood, and Fire-Related Disaster as shown in Table 3. Multi-hazard apps (27 apps) cover different types of disasters. For example, ‘Disaster Alert’ app covers 18 types of disasters, ranging from earthquakes and floods to wildfires. Earthquake apps (15 apps), like ‘Earthquakes Today’, exclusively cover earthquakes, while others, such as ‘Safety Tips’, also encompass volcanic activity and tsunamis. Wind disaster apps (8 apps) primarily address three kinds of wind events: hurricanes, cyclones, and typhoons. The last two single types of disasters were floods (4 apps) and fires (1 app), with the latter covering bushfires, forest fires, or other fire-related disasters (refer to Table 3 to view the number of apps in each category). The detailed lists of these apps can be found in Appendix Tables 10 and11 for Playstore and Table 12 for AppStore.
Table 2: Apps by Disaster Type
@@
Multi-Hazard   22
Earthquakes   12
@@
Analysis 3 – Geographic Coverage: Geographic locations supported in disaster apps can be categorised as international and local. The local apps can have sub-categories of region-specific, national and territorial. Intentional apps (22 apps) can be accessed from any location and cover all countries. Different local apps can be accessed from only those geographical areas. Region-specific apps (6 apps), cover regions such as the USA and Canada or the USA and Europe (see Appendix Table 9). Some regions may also extend to various basins or oceans, particularly for apps focusing on wind-related disasters. National apps (20 apps) exclusively target a single country. Lastly, territorial apps (8 apps), such as ‘VicEmergency’, exclusively cover disasters within a specific sub-national area, such as Victoria, Australia, while ‘SD Emergency’ focuses solely on disasters in San Diego, California, USA. The detailed lists of these apps along with their geographic coverage can be found in Appendix Tables 10 and11 for Playstore and Table 12 for AppStore.
Table 3: Apps by Geographic Coverage
@@
International   19
Region-Specific   6
National   19
Territorial   5
Analysis 4 – Popularity and Quality: On the Google Play Store, the rating of an app is given in a numerical range from 1 to 5, and the number of downloads is expressed within specific ranges, such as 10K+, 500K+, 1M+, and 5M+. Meanwhile, on the App Store, the app rating follows the same scale, ranging from 1 to 5. However, the App Store provides precise figures without rounding to the nearest thousand or million for the number of downloads. The specific number of downloads for each of the apps we explored can be found in Appendix Tables 10, 11 and 12.
Table 4: Comparison of Total Downloads and App Ratings on Play Store and App Store
Variable Statistics Play Store App Store
@@
Of the 45 apps in this study, some of are published in both the Play Store and App Store. The number of disaster-related apps in the Play Store is 41, while in the App Store, there are 16. On average and median (see Table 4), the rating value in the Play Store surpasses that of the App Store. The average and median ratings on the Play Store are 3.88 and 4.1, respectively, whereas on the App Store, they stand at 3.68 and 3.75. The app, ‘BD999’, with the highest rating, is also on the Play Store, which received a perfect rating of 5. Meanwhile, the app, ‘SES Assistance QLD’, with the lowest rating, is on the App Store, which received a rating of 1.7.
The number of downloads for all disaster-related apps in the Play Store exceeded 22 million, while in the App Store, it is over 13,000. Statistically, the average and median number of downloads on the Play Store are over 559,000 and 100,000, respectively, while on the App Store, they are 840 and 435. What’s interesting here is that the app with the lowest number of downloads on the Play Store has almost the same value as the one with the most on the App Store. However, the significant difference in the number of downloads in these two stores is reasonable, considering the higher percentage of Android users compared to iOS, with Android users accounting for 70.11% and iOS users for 29.19% [35]. The app ‘NINA - Die Warn-App des BBK’ has the most downloads is in the Play Store with more than 5 million downloads. The app ‘SES Assistance QLD’ has the fewest downloads is in the App Store, only downloaded 6 times.
Analysis 5 – Monetization Strategies: Developers can earn money from the apps they create through ads, in-app purchases, and/or purchases made during the download process. More than half (24 apps) of disaster apps were free, while the other half used a combination of pricing strategies. As shown in Table 5, 10 apps on the Play Store generate revenue by incorporating ads within their apps, offering premium feature packages for purchase, or providing options for ad removal, either individually or in combination. Additionally, there are 3 apps that require users to pay when downloading them. Meanwhile, the sole method for generating revenue in the App Store is offering in-app purchase services. The monetization strategy for each of the apps we explored can be found in Appendix Tables 10, 11 and 12.
Table 5: Pricing Schema of Natural Disaster App
Pricing Schema Play App
Store Store
Free 24 12
@@
Free with In-App Purchase 2 2

### v2

@@ 3.2.2 A Dataset of Disaster App Reviews
Collecting Reviews: To answer RQ2 and RQ3, we needed user reviews of the 44 disaster apps. To gather user reviews of 44 disaster apps from the Google Play Store and the App Store, we utilized an app-store-scraper  [^6] https://github.com/cowboy-bebug/app-store-scraper and google-play-scraper  [^7] https://github.com/JoMingyu/google-play-scraper. These tools require the app’s ID, name, and the specific country of the store. The app’s name is visible upon selection, whereas its ID and the relevant country code are identifiable through its URL. We collected 37,272 user reviews from 44 disaster apps.
Translation: The collected reviews encompassed a diverse linguistic range, particularly for applications originating from non-English-speaking countries. Examples include DEC112 2.0 (German), Zachranka (Czech), and Yurekuru Call (Japanese). Among the 44 apps analyzed, 9 contained reviews in languages other than English, amounting to 2,793 reviews that required translation. To address this, we employed the OPUS-MT Zero-Shot Translation [Tiedemann2023] model to translate non-English reviews into English. We selected OPUS-MT because it represents one of the most comprehensive collections of machine translation models. The OPUS-MT repository comprises more than 2,300 models, supporting 4,560 language directions across 294 languages [Tiedemann2023]. Furthermore, OPUS-MT achieves competitive translation quality, with BLEU scores comparable to large-scale systems such as NLLB [nllbteam2022languageleftbehindscaling]
@@
Exclusions: The next step was to exclude brief reviews comprising fewer than four words. The reason behind this exclusion is short reviews offering minimal commentary such as ‘good’, ‘not accurate’, and ‘works well’. Thus, this type of reviews provide insufficient information for meaningful analysis and can’t be used to generate a topic for RQ2. In addition, reviews without a rating were removed, as they could not be used in the sentiment analysis. After this filtering, the dataset was reduced to 28,161 reviews, providing a solid foundation for the analysis of RQ2 and RQ3.
3.3 Methodology Approach
3.3.1 RQ1: Key features of disaster apps
Analysis 1- App Features: To identify the available app features, we conducted a hands-on evaluation by downloading and using each application directly. This approach allowed us to explore the full range of functionalities beyond what was described in the app listings or shown in image previews. During this process, we noted each feature observed through actual use and interaction with the app interface. The collected features were then categorized into three main types, aligned with the four stages of the disaster lifecycle: preparation, response, recovery, and mitigation [04Unit4d19_online].
Analysis 2- Disaster Type Coverage: In addition to categorizing apps based on their lifecycle phase, we also examined the types of disasters each app addresses. Apps were classified into two groups: single-hazard apps and multi-hazard apps. Single-hazard apps have features that focus on only one type of disaster, such as an earthquake, while multi-hazard apps cover multiple types of disasters.
Analysis 3- Geographic Coverage: Apps can also be classified based on the geographic scope of their coverage. We categorized the apps into four levels: local, national, regional or territorial, and global. Local apps provide coverage for areas smaller than a nation, such as a state, province, or city. National apps cover a single nation. Regional or territorial apps cover multiple nations or areas, such as the Pacific regions. Global apps provide coverage all nations.
Analysis 4- Popularity and Quality Analysis: The popularity of an app can be gauged by its number of downloads, while user ratings offer insights into its perceived quality. These data can be obtained from the homepage of each app on the Google Play Store and the App Store. For example, Figure 2 shows that the ‘Hurricane Tracker Pro’ app has been downloaded 5K+ and has a rate of 4.7.
Figure 2: App pricing scheme, number of downloads, and rating
Analysis 5- Monetization Strategies: Not all disaster apps are freely available. To determine whether an app is free, we examined the homepage of each app, which typically indicates if the app contains advertising or offers in-app purchases (payments within the app to access certain features or remove advertisements). For example, as shown in Figure 2, users must pay \$2.99 to download the ‘Hurricane Tracker Pro’ app. In contrast, the ‘Wind Map Hurricane Tracker 3D’ app, also shown in Figure 2, is free to download but contains advertising and offers in-app subscriptions.
3.3.2 RQ2: User Discussions in App Reviews
To understand the topics discussed by users of disaster apps, we applied BERTopic [grootendorst2022bertopic] to our review dataset. We then performed cluster merging, as clustering algorithms rely on distance metrics and do not inherently recognize when separate clusters represent the same topic. This allowed us to combine similar clusters. After merging, we refined the clusters by removing outliers within each cluster. The overall topic modeling process is illustrated in Figure 3, and the details of the approach are described below.
Figure 3: Topic Modeling Process
BERTopic: In this research, BERTopic was utilized for topic modeling, as it is a state-of-the-art model for clustering tasks. BERTopic converts documents into vectors using the Sentence-BERT (SBERT) [reimers2019sentencebertsentenceembeddingsusing] framework. SBERT is utilized for document embedding because it can retain semantic information. BERTopic provides two options for sentence embedding: all-MiniLM-L6-v2888https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2 for English sentences and paraphrase-multilingual-MiniLM-L12-v2999https://huggingface.co/sentence-transformers/paraphrase-multilingual-MiniLM-L12-v2 for multilingual sentences. The first option was chosen because we translated all non-English reviews into English, so we don’t need a multilingual model for sentence embedding. Subsequently, the dimensionality of the embedding vector is reduced using UMAP (Uniform Manifold Approximation and Projection) [mcinnes2020umap].
UMAP was selected for dimension reduction in BERTopic for its superior embedding stability and significantly faster runtime compared to other methods such as T-SNE [JMLR:v9:vandermaaten08a]. Once reduced, the embedding vector is clustered using HDBSCAN (Hierarchical density-based clustering) [mcinnes2017hdbscan]. HDBSCAN was chosen because this model can prevent unrelated documents (outliers) from being assigned to any cluster and is more robust in parameter selection. HDBSCAN has parameters that need to be specified, among which is min_topic_size [grootendorst2023parametertuning]. This parameter ensures that each cluster has at least min_topic_size members. If there are numerous points in close proximity but fewer than the min_topic_size, these points will either join another cluster or become outliers. In this study, the min_topic_size used in this clustering was 50.
Cluster Merging and Naming: After the model training is complete, BERTopic will provide results in the form of topic IDs, topic names, keywords, and the number of members in each topic. To ensure the accuracy of the topic naming, a manual analysis is conducted to rename each topic. This analysis involves examining the keywords of each topic, as well as checking 50 data points that are closest in distance to the center of each cluster. After all clusters are appropriately named, similar clusters are merged, for example, clusters named “app notification” and “app alert” are considered the same topic.
Cluster Refinement: In the topic modeling process, each identified topic is refined by removing reviews that do not align well with the topic’s main theme. This is done by using BERTopic to re-cluster the reviews within a topic into two groups: outliers and non-outliers. Reviews classified as outliers are discarded, ensuring that only the relevant reviews are retained as the final result of the topic modeling process. If a topic contains no outliers, the number of reviews remains unchanged. Finally, every user review is assigned to a single cluster since BERTopic is not soft clustering method. If a review contains multiple topics, we assign it to the most dominant topic.
3.3.3 RQ3: Challenges in Disaster Apps
Figure 4: Sentiment Analysis Process
Sentiment Analysis: This research question examines which topics have the highest percentage of negative reviews. This information can help developers identify dominant user complaints and prioritize their resolution. Ratings are commonly used to indicate user satisfaction with an app. However, Text-Rating Inconsistency (TRI) can occur when users provide ratings carelessly [BinFu2013]. For instance, a user may give a positive rating immediately after downloading the app without actually using it, or a user may submit numerous complaints yet assign a neutral rating, such as a score of three. Relying solely on ratings may therefore fail to accurately reflect user sentiment. TRI can also result in mislabeled data, which reduces the reliability of supervised learning approaches. To address this issue, unsupervised methods that do not depend on label accuracy can be applied, with VADER [hutto2014vader] being a particularly effective option.
The process of sentiment analysis is shown in Figure 4. The outcome of VADER includes four types of sentiment scores: negative scores, neutral scores, positive scores, and compound scores. VADER considers that each sentence may contain both negative and positive elements. For example, consider the review “Currently, I like the app very much, but you should be able to change the language within the app.” This review contains a positive element: “I like the app very much” but also a negative element: “you should be able to change the language”.
VADER provides a combined score of these elements, known as a compound score, in addition to the individual sentiment scores. This compound score, ranging from -1 to 1, will be used as the result of sentiment analysis in this study. We will classify sentiment as positive (score >= 0.05), neutral (0.05 > score > -0.05), or negative (score <= -0.05), following similar approaches used in prior work [Chembakottu2023].
Topic with Lowest Rating and Sentiment Score: In this analysis, we examine the user rating and sentiment score for each topic. Next, we identify which topics have the lowest score, enabling developers to prioritize these topics for improvement.
@@ 4.1 RQ1: Key features of disaster apps
Analysis 1 — App Features: Each disaster support app provides one or more features related to natural disasters. The following section explains the key features we categorized according to the stages of a disaster event: pre-disaster (preparation), during-disaster (response), and post-disaster (recovery). The number of apps offering each feature type is shown in Table 1.
Table 1: App feature categorization and number of apps
@@
      Early Warning Alerts: This feature enables users to receive immediate information when a disaster is about to occur, allowing them to take precautionary actions to protect themselves. Apps offering this feature are commonly developed in Japan, particularly those related to earthquakes. A good example is the Yukureku Call app (see Figure LABEL:fig_ew).
@@
      Disaster Forecasting: This feature supports the prediction of potential disasters at specific times and places. Early warning alerts are sent moments before a disaster and are almost certain to occur. In contrast, forecasting provides a longer lead time, giving users more time to prepare. However, its accuracy is lower than that of early warning systems. This feature is common in apps that monitor wind movements, as typhoons and hurricanes can be predicted more reliably than other disasters. An example is the Seastorm Hurricane Tracker app (see Figure LABEL:fig_forecast).
@@
      Real Time Alerts: This alert appears as some form of pop-up, informing the user of a disaster happening in the area according to their radius settings or specified area of interest. This feature differs from an early warning alert because notifications arrive during a disaster, not beforehand. For instance, if a user sets a threshold for earthquakes measuring 3 on the Richter scale and above, they will receive a notification when such an event occurs. Such an example is provided in ‘QuakeFeed Earthquake Tracker’ app (see Figure LABEL:fig_at_disaster_1).
@@
Table 2: Disaster Type Distribution of Mobile Apps (The Google Play Store and The App Store)
@@
Multi-Hazard   26
Earthquakes   16
@@
Table 3: Geographic Coverage Distribution of Mobile Apps (The Google Play Store and The App Store)
@@
Global   22
Region / Territorial   6
National   20
Local   8
Analysis 2 — Disaster Type Coverage: Based on the types of disasters covered, the apps can be divided into two categories: multi-hazard apps and single-hazard apps. Multi-hazard apps cover various types of disasters. For example, the Disaster Alert app covers 18 types of disasters. In contrast, single-hazard apps focus on only one type of disaster. From our review, we found that some apps focus solely on earthquakes, some on wind-related disasters, others on floods, and some on fire-related disasters. The distribution of app types based on the disasters they cover is shown in Table 2. Detailed information on which disasters each app covers can be found in Table A2 and Table LABEL:tbl:list_of_play_store_app.
Analysis 3 – Geographic Coverage: Geographic locations supported in disaster apps can be categorized as global and non-global. The non-global category includes regional, national, and local apps. Global apps can be accessed from any location and cover all countries. In contrast, non-global apps are limited to specific geographical areas. Regional apps cover broad areas such as the USA and Canada or the USA and Europe (see Appendix Table A1). Some regional apps also extend across oceanic basins, especially those focusing on wind-related disasters. National apps target a single country only. Territorial apps, such as ‘VicEmergency’, focus on a sub-national area like Victoria, Australia, while ‘SD Emergency’ covers only San Diego, California, USA. The distribution of apps based on coverage is shown in Table 3. Detailed lists of these apps and their geographic coverage can be found in Appendix Table LABEL:tbl:list_of_play_store_app for the Google Play Store and Table A2 for the App Store.
Analysis 4 – Popularity and Quality: On the Google Play Store, the rating of an app is given in a numerical range from 1 to 5, and the number of downloads is expressed within specific ranges, such as 10K+, 500K+, 1M+, and 5M+. Meanwhile, on the App Store, the app rating follows the same scale, ranging from 1 to 5. However, the App Store provides precise figures without rounding to the nearest thousand or million for the number of downloads. The specific number of downloads for each of the apps we explored can be found in Appendix Table A2 and Table LABEL:tbl:list_of_play_store_app.
Table 4: Comparison of Total Downloads and App Ratings
Variable Statistic The Google Play Store The App Store
@@
Table 5: Pricing Schema of Natural Disaster Apps
Pricing Schema
The Google
Play Store
The
App Store
Free 22 12
@@
Free with In-App Purchase 3 3
@@
Of the 44 apps in this study, some of are published in both the Google Play Store and the App Store. The number of disaster-related apps in the Google Play Store is 40, while in the App Store, there are 16. On average and median (see Table 4), the rating value in the Google Play Store surpasses that of the App Store. The average and median ratings on the Google Play Store are 3.88 and 4.1, respectively, whereas on the App Store, they stand at 3.68 and 3.75. The app, ‘BD999’, with the highest rating, is also on the Google Play Store, which received a perfect rating of 5. Meanwhile, the app, ‘SES Assistance QLD’, with the lowest rating, is on the App Store, which received a rating of 1.7.
The number of downloads for all disaster-related apps in the Google Play Store exceeded 22 million, while in the App Store, it is over 13,000. Statistically, the average and median number of downloads on the Google Play Store are over 559,000 and 100,000, respectively, while on the App Store, they are 840 and 435. What’s interesting here is that the app with the lowest number of downloads on the Google Play Store has almost the same value as the one with the most on the App Store. However, the significant difference in the number of downloads in these two stores is reasonable, considering the higher percentage of Android users compared to iOS, with Android users accounting for 70.11% and iOS users for 29.19% [StatistaMobileOS]. The app ‘NINA - Die Warn-App des BBK’ has the most downloads is in the Google Play Store with more than 5 million downloads. The app ‘SES Assistance QLD’ has the fewest downloads is in the App Store, only downloaded 6 times.
Analysis 5 – Monetization Strategies: Developers can earn revenue from the apps they create through advertisements, in-app purchases, paid downloads, or a combination of these methods. More than half (22 apps) of the disaster apps were free, while the remaining apps adopted various monetization strategies. As shown in Table 5, five apps on the Google Play Store generate income through in-app advertisements. Three apps offer premium feature packages or provide options to remove ads. Eight apps include both advertisements and premium features available for purchase. One app requires payment upon download, while another requires payment and also offers an upgrade option. Meanwhile, the primary monetization approach in the App Store involves offering in-app purchase services, either as free apps with optional purchases or as paid downloads. The detailed monetization strategy for each app is provided in Appendix Table A2 and Table LABEL:tbl:list_of_play_store_app.